# On the second Bohr radius for vector valued pluriharmonic functions

### v1

@@ Abstract.
In this paper, we introduce the notion of the second Bohr radius for vector valued pluriharmonic functions on complete Reinhardt domains in $\mathbb{C}^{n}$. This investigation is motivated by the work of Lev Aizenberg [Proc. Amer. Math. Soc. 128 (2000), 1147 –1155], where the corresponding problem was studied for complex valued holomorphic functions. We show that the second Bohr radius constant for pluriharmonic functions is strictly positive under suitable condition. In addition, we obtain its asymptotic behavior in the finite-dimensional settings using invariants from local Banach space theory. Asymptotic estimates for this constant are obtained on both convex and non-convex complete Reinhardt domains. Our results also apply to a broad class of Banach sequence spaces, including symmetric and convex Banach spaces. The framework developed here also includes the second Bohr radius problem for vector valued holomorphic functions. As an application of our results, we derive several consequences that extend known results in the scalar valued setting as well as existing results in the literature.
@@ 1. Introduction and the main results
The paper is organized as follows. In Section $2$, we find the condition and show that second Bohr radii constants $B_{\lambda}(\Omega,p,U)$ and $L_{\lambda}(\Omega,p,U)$ are strictly positive. Section $3$ is devoted to the asymptotic study of these constants in the case where $X$ or $\mathcal{B}(\mathcal{H})$ is finite dimensional.
@@ 2. Characterization of non zero second Bohr radii constants
Before stating our result, we introduce the following standard notation. For bounded Reinhardt domains $\Omega_{1},\Omega_{2}\subset\mathbb{C}^{n}$, let $S(\Omega_{1},\Omega_{2}):=\inf\left\{s>0:\Omega_{1}\subset s\,\Omega_{2}\right\}$. A Banach space $Z$ satisfying $\ell_{1}\subset Z\subset c_{0}$ (with normal inclusions) is called a Banach sequence space if the canonical unit vectors $e_{k}$} form a $1$-unconditional basis of $Z$ (see Section $3$ for the definition of unconditional basis). Recall that if $Z$ and $W$ are Banach sequence spaces then $S(B_{Z},B_{W})=\left\lVert Id:Z\rightarrow W\right\rVert$.
@@ 3. Asymptotic estimates of the second Bohr radii constants in finite dimensional settings
In particular, when $n=2$ this yields the bounds $0.23570226\leq K_{2}(\mathbb{D}^{2})\leq 1/3=0.33333$. After nearly three decades, these estimates have been improved in recent years. Specifically, Knese showed in $2025$ that $K_{2}(\mathbb{D}^{2})>0.3006$, while Baran, Pikul, Woerdeman, and Wojtylak proved in $2026$ that $K_{2}(\mathbb{D}^{2})<0.3177$. Taken together, these results confirm in particular that $K_{n}(\mathbb{D}^{n})<1/3$ for $n\geq 2$.
@@
$\frac{1}{3}\,\left(\frac{1}{n}\right)^{\frac{1}{2}+\frac{1}{\max\{q,2\}}}\leq B(B_{\ell^{n}_{q}},1)\leq 4\,\left(\frac{\log\,n}{n}\right)^{\frac{1}{2}+\frac{1}{\max\{q,2\}}}.$
@@
Let $Z=(\mathbb{C}^{n},||.||)$ be a Banach space such that $\chi(\{e_{k}\}^{n}_{k=1})=1$. Let $X=\mathcal{B}(\mathcal{H})$ be a finite dimensional complex Banach space and $\lambda>1$. Then for $1\leq q\leq\infty$,
$L_{\lambda}(B_{Z},p,X)\geq\begin{cases}E_{1}(X)\,\frac{1}{s\,t\,n^{\frac{1}{q}}}\max\left\{\frac{1}{\sqrt{n}\,\left\lVert Id:\ell^{n}_{2}\rightarrow Z\right\rVert},\,\frac{1}{e\,\left\lVert Id:Z\rightarrow\ell^{n}_{1}\right\rVert}\right\}\,\left(\frac{\lambda^{p}-1}{2\lambda^{p}-1}\right)^{\frac{1}{p}}&\text{for $p=1$},\\[5.69054pt] E_{3}\,\frac{1}{s}\,\left\lVert I:\ell^{n}_{2}\rightarrow Z\right\rVert^{-\frac{2}{p}}\,\left(\frac{\lambda^{p}-1}{2\lambda^{p}-1}\right)^{\frac{1}{p}}\,&\text{for $p\geq 2$},\\[5.69054pt] \frac{E_{2}(X)}{s\left(tn^{\frac{1}{q}}\right)^{\frac{(2-p)}{p}}}\max\left\{\frac{1}{(\sqrt{n})^{1-\theta}\left\lVert Id:\ell^{n}_{2}\rightarrow Z\right\rVert},\frac{\left\lVert Id:\ell^{n}_{2}\rightarrow Z\right\rVert^{-\theta}}{e^{(1-\theta)}\left\lVert Id:Z\rightarrow\ell^{n}_{1}\right\rVert^{1-\theta}}\right\}\left(\frac{\lambda^{p}-1}{2\lambda^{p}-1}\right)^{\frac{1}{p}}\hskip-8.53581pt&\text{for $1<p<2$},\end{cases}$
where $\theta=(2(p-1))/p$, $s:=\left\lVert Id:Z\rightarrow\ell^{n}_{q}\right\rVert$, and $t:=\left\lVert Id:\ell^{n}_{q}\rightarrow Z\right\rVert$. Here, $E_{1}(X)$, $E_{2}(X)$, and $E_{3}$ are positive constants: $E_{1}(X)$ depends only on $X$, $E_{2}(X)$ depends on both $X$ and $p$, while $E_{3}$ is independent of $X$ and depends only on $p$. Furthermore,
$L_{\lambda}(B_{Z},p,X)\leq d\,\lambda^{\frac{1}{\log\,n}}\,b_{[\log\,n]}(B_{Z})\,\,\frac{\left\lVert Id:Z\rightarrow\ell^{n}_{2}\right\rVert}{n^{\frac{1}{p}}}$
@@
The correspondence preceding Theorem 3.1 allows Theorem 3.1 to be extended to bounded simply connected convex complete Reinhardt domains. Moreover, combining Theorem 3.1 and Lemma 2.1 yields the asymptotic estimates for all bounded simply connected complete Reinhardt domains.
@@ 3.1. Basic definitions and notations
$Cot(W):=\inf\{2\leq t\leq\infty:W\,\mbox{has cotype}\,t\}.$
It is well known that every Banach space has cotype $\infty$. We adopt the convention $1/Cot(W)=0$ whenever $Cot(W)=\infty$.
@@ 3.2. Second Bohr radii for homogeneous plynomials
We want to first obtain the desired lower bound of $R^{m}_{\lambda}(B_{Z},p,X)$ when $X$ is finite dimensional. For all $m\geq 1$, let $P(z)=\sum_{|\alpha|=m}a_{\alpha}\,z^{\alpha}+\sum_{|\beta|=m}b^{*}_{\alpha}\,\bar{z}^{\alpha}\in\mathcal{BPH}(^{m}B_{Z},X)$. Then for $r\in[0,1)$,
(3.3) $\sum_{|\alpha|=m}\sup_{z\in r\,B_{Z}}(\left\lVert a_{\alpha}\right\rVert^{p}+\left\lVert b_{\alpha}\right\rVert^{p})|z|^{p\alpha}\leq r^{pm}\,\left\lVert Id:Z\rightarrow\ell^{n}_{q}\right\rVert^{pm}\,\sum_{|\alpha|=m}(\left\lVert a_{\alpha}\right\rVert^{p}+\left\lVert b_{\alpha}\right\rVert^{p})\,.$
We now need to estimate the term $\sum_{|\alpha|=m}(\left\lVert a_{\alpha}\right\rVert^{p}+\left\lVert b_{\alpha}\right\rVert^{p})$. To do so, we consider several cases depending on the values of $p$. More precisely, the proof is presented in three separate cases, depending on the values of $p$: $p=1$, $p\geq 2$, and $1<p<2$. Let $H$ and $G$ be the functions as defined in the proof of Theorem 2.1. Set $s:=\left\lVert Id:Z\rightarrow\ell^{n}_{q}\right\rVert$ and $t:=\left\lVert Id:\ell^{n}_{q}\rightarrow Z\right\rVert$.
@@
(3.8) $\chi_{M}(\mathcal{P}(^{m}Z))\leq n^{\frac{m}{2}}\,\left\lVert Id:\ell^{n}_{2}\rightarrow Z\right\rVert^{m}\,\,\mbox{and}\,\,\chi_{M}(\mathcal{P}(^{m}Z))\leq\frac{m^{m}}{m!}\,\left\lVert Id:Z\rightarrow\ell^{n}_{1}\right\rVert^{m},$
@@
$L^{m}_{\lambda}(B_{Z},p,X)\geq E_{1}(X)\,\,\frac{\lambda^{\frac{1}{m}}}{s\,t\,n^{\frac{1}{q}}}\,\max\left\{\frac{1}{\sqrt{n}\,\left\lVert Id:\ell^{n}_{2}\rightarrow Z\right\rVert},\,\frac{1}{\left(\frac{m^{m}}{m!}\right)^{1/m}\,\left\lVert Id:Z\rightarrow\ell^{n}_{1}\right\rVert}\right\},$
@@
$\left(\sum_{|\alpha|=m}\left|\phi\left(\frac{a_{\alpha}+b_{\alpha}}{2}\right)\right|^{2}\right)^{\frac{1}{2}}\leq\left\lVert Id:\ell^{n}_{2}\rightarrow Z\right\rVert^{m}\left\lVert H\right\rVert_{B_{Z},X}$
@@
(3.9) $\sum_{|\alpha|=m}\left\lVert\left(\frac{a_{\alpha}+b_{\alpha}}{2}\right)\right\rVert^{p}\leq 2^{p-2}\,\left\lVert Id:\ell^{n}_{2}\rightarrow Z\right\rVert^{2m}\left\lVert H\right\rVert^{2}_{B_{Z},X}.$
@@
(3.10) $\sum_{|\alpha|=m}\left\lVert\left(\frac{a_{\alpha}-b_{\alpha}}{2}\right)\right\rVert^{p}\leq 2^{p-2}\,\left\lVert Id:\ell^{n}_{2}\rightarrow Z\right\rVert^{2m}\left\lVert G\right\rVert^{2}_{B_{Z},X}.$
@@
$\left(\sum_{|\alpha|=m}\left\lVert a_{\alpha}\right\rVert^{p}\right)^{\frac{1}{p}}\leq 2\left\lVert Id:\ell^{n}_{2}\rightarrow Z\right\rVert^{\frac{2m}{p}}2^{\frac{2}{p}}.$
@@
$\sum_{|\alpha|=m}\sup_{z\in r\,B_{Z}}(\left\lVert a_{\alpha}\right\rVert^{p}+\left\lVert b_{\alpha}\right\rVert^{p})|z|^{p\alpha}\leq 2^{p+3}\,r^{pm}\,s^{pm}\,\left\lVert Id:\ell^{n}_{2}\rightarrow Z\right\rVert^{2m}\leq\lambda^{p}$
@@
$r\leq\frac{\lambda^{\frac{1}{m}}}{2^{\frac{p+3}{pm}}\,s\,\left\lVert Id:\ell^{n}_{2}\rightarrow Z\right\rVert^{\frac{2}{p}}}.$
@@
$L^{m}_{\lambda}(B_{Z},p,X)\geq\frac{\lambda^{\frac{1}{m}}}{2^{\frac{p+3}{pm}}\,s\,\left\lVert Id:\ell^{n}_{2}\rightarrow Z\right\rVert^{\frac{2}{p}}}.$
@@
$\sum_{|\alpha|=m}\left\lVert\left(\frac{a_{\alpha}+b_{\alpha}}{2}\right)\right\rVert^{p}\leq 2^{p}\,(\pi_{1}(I_{X}))^{2-p}\,\left(t\,n^{\frac{1}{q}}\right)^{(2-p)m}(\chi_{M}(\mathcal{P}(^{m}Z)))^{2-p}\left\lVert Id:\ell^{n}_{2}\rightarrow Z\right\rVert^{2(p-1)m}.$

### v2

# arXiv:2601.12608v2[math.CV] 27 Feb 2026

@@ Abstract.
In this paper, we introduce the notion of the second Bohr radius for vector valued pluriharmonic functions on complete Reinhardt domains in $\mathbb{C}^{n}$. This investigation is motivated by the work of Lev Aizenberg [Proc. Amer. Math. Soc. 128 (2000), 1147 –1155], where the corresponding problem was studied for complex valued holomorphic functions. We show that the second Bohr radius constant for pluriharmonic functions is strictly positive under suitable condition. In addition, we obtain its asymptotic behavior in both the finite- and infinite-dimensional settings using invariants from local Banach space theory. Asymptotic estimates for this constant are obtained on both convex and non-convex complete Reinhardt domains. Our results also apply to a broad class of Banach sequence spaces, including symmetric and convex Banach spaces. The framework developed here also includes the second Bohr radius problem for vector valued holomorphic functions. As an application of our results, we derive several consequences that extend known results in the scalar valued setting as well as existing results in the literature.
@@ 1. Introduction
The paper is organized as follows. In Section $2$, we establish conditions under which the second Bohr radii constants $B_{\lambda}(\Omega,p,U)$ and $L_{\lambda}(\Omega,p,U)$ are strictly positive. Section $3$ is devoted to the asymptotic study of these constants using invariants from local Banach space theory in the case where $X$ or $\mathcal{B}(\mathcal{H})$ is finite-dimensional, while the infinite-dimensional setting is treated in Section $4$. Section $5$ presents several applications of our results to Banach sequence spaces.
@@ 2. Characterization of non zero second Bohr radii constants
Before stating our result, we introduce the following standard notation. For bounded Reinhardt domains $\Omega_{1},\Omega_{2}\subset\mathbb{C}^{n}$, let $S(\Omega_{1},\Omega_{2}):=\inf\left\{s>0:\Omega_{1}\subset s\,\Omega_{2}\right\}$. A Banach space $Z$ satisfying $\ell_{1}\subset Z\subset c_{0}$ (with normal inclusions) is called a Banach sequence space if the canonical unit vectors $e_{k}$} form a $1$-unconditional basis of $Z$ (see Section $3$ for the definition of unconditional basis). Recall that if $Z$ and $W$ are Banach sequence spaces then $S(B_{Z},B_{W})=\left\lVert\mathrm{Id}:Z\rightarrow W\right\rVert$.
@@
Proof.
We first prove the right hand inequality. Let $f\in\mathcal{PH}(\Omega_{2},\mathcal{B}(\mathcal{H}))$ be of the form (1.2) with $\|f\|_{\Omega_{2}}\leq 1$.
Let $\delta_{1}>0$. By the definition of $S(\Omega_{1},\Omega_{2})$,
$z\in\Omega_{1}\quad\Rightarrow\quad\frac{z}{S(\Omega_{1},\Omega_{2})+\delta_{1}}\in\Omega_{2}.$
Define
$g(z)=f\!\left(\frac{z}{S(\Omega_{1},\Omega_{2})+\delta_{1}}\right).$
Then $g\in\mathcal{PH}(\Omega_{1},\mathcal{B}(\mathcal{H}))$ and $\|g\|_{\Omega_{1}}\leq 1$.
Now, for fix $0<\varepsilon<L_{\lambda}(\Omega_{1},p,U)$, let $r=L_{\lambda}(\Omega_{1},p,U)-\varepsilon$. Since $r<L_{\lambda}(\Omega_{1},p,U)$, we have
$\sum_{m=0}^{\infty}\sum_{|\alpha|=m}\sup_{z\in r\Omega_{1}}(\|U(a_{\alpha})\|^{p}+\|U(b_{\alpha})\|^{p})\left|\left(\frac{z}{S(\Omega_{1},\Omega_{2})+\delta_{1}}\right)^{\alpha}\right|^{p}\leq\lambda^{p}.$
Hence
(2.1) $\sum_{m=0}^{\infty}\sum_{|\alpha|=m}\sup_{z\in\frac{r}{S(\Omega_{1},\Omega_{2})+\delta_{1}}\Omega_{1}}(\|U(a_{\alpha})\|^{p}+\|U(b_{\alpha})\|^{p})|z^{\alpha}|^{p}\leq\lambda^{p}.$
Now let $\delta_{2}>0$. Using the definition of $S(\Omega_{2},\Omega_{1})$, we have
$z\in\Omega_{2}\quad\Rightarrow\quad\frac{z}{S(\Omega_{2},\Omega_{1})+\delta_{2}}\in\Omega_{1}.$
Thus, from (2.1), we obtain
$\sum_{m=0}^{\infty}\sum_{|\alpha|=m}\sup_{z\in\frac{r}{(S(\Omega_{1},\Omega_{2})+\delta_{1})(S(\Omega_{2},\Omega_{1})+\delta_{2})}\Omega_{2}}(\|U(a_{\alpha})\|^{p}+\|U(b_{\alpha})\|^{p})|z^{\alpha}|^{p}\leq\lambda^{p}.$
By definition of $L_{\lambda}(\Omega_{2},p,U)$,
$\frac{r}{(S(\Omega_{1},\Omega_{2})+\delta_{1})(S(\Omega_{2},\Omega_{1})+\delta_{2})}\leq L_{\lambda}(\Omega_{2},p,U).$
Letting $\delta_{1},\delta_{2}\to 0$ gives
$r\leq S(\Omega_{1},\Omega_{2})S(\Omega_{2},\Omega_{1})L_{\lambda}(\Omega_{2},p,U).$
Since $\varepsilon$ is arbitrary, we deduce that
$L_{\lambda}(\Omega_{1},p,U)\leq S(\Omega_{1},\Omega_{2})S(\Omega_{2},\Omega_{1})L_{\lambda}(\Omega_{2},p,U).$
Interchanging $\Omega_{1}$ and $\Omega_{2}$ in the preceding inequality gives the left hand inequality. ∎
The above lemma can be viewed as an analogue of [20, Lemma 3.1(1)], where the authors considered only scalar valued holomorphic functions, with $U$ being the identity operator on $\mathbb{C}$ and the power $p=1$. The proof follows the same pattern as that of [20, Lemma 2.5]. For the sake of completeness, we include the proof here. The main difference lies in the fact that we work with vector valued pluriharmonic functions associated with a bounded operator and an arbitrary power $p$.
@@ 3. Asymptotic estimates of the second Bohr radii constants in finite dimensional settings
In particular, when $n=2$ this yields the bounds $0.23570226\leq K_{2}(\mathbb{D}^{2})\leq 1/3=0.33333$. After nearly three decades, these estimates have been improved in recent years. Specifically, Knese [31] showed in $2025$ that $K_{2}(\mathbb{D}^{2})>0.3006$, while Baran, Pikul, Woerdeman, and Wojtylak [5] proved in $2026$ that $K_{2}(\mathbb{D}^{2})<0.3177$. Taken together, these results confirm in particular that $K_{n}(\mathbb{D}^{n})<1/3$ for $n\geq 2$.
@@
(3.1) $\frac{1}{3}\,\left(\frac{1}{n}\right)^{\frac{1}{2}+\frac{1}{\max\{q,2\}}}\leq B(B_{\ell^{n}_{q}},1)\leq 4\,\left(\frac{\log\,n}{n}\right)^{\frac{1}{2}+\frac{1}{\max\{q,2\}}}.$
@@
Let $Z=(\mathbb{C}^{n},||.||)$ be a Banach space such that $\chi(\{e_{k}\}^{n}_{k=1})=1$. Let $X=\mathcal{B}(\mathcal{H})$ be a finite dimensional complex Banach space and $\lambda>1$. Then for $1\leq p<\infty$, $1\leq q\leq\infty$,
$L_{\lambda}(B_{Z},p,X)\geq\begin{cases}E_{1}(X)\,\frac{1}{s\,t\,n^{\frac{1}{q}}}\max\left\{\frac{1}{\sqrt{n}\,\left\lVert\mathrm{Id}:\ell^{n}_{2}\rightarrow Z\right\rVert},\,\frac{1}{e\,\left\lVert\mathrm{Id}:Z\rightarrow\ell^{n}_{1}\right\rVert}\right\}\,\left(\frac{\lambda^{p}-1}{2\lambda^{p}-1}\right)^{\frac{1}{p}}&\text{for $p=1$},\\[5.69054pt] E_{3}\,\frac{1}{s}\,\left\lVert\mathrm{Id}:\ell^{n}_{2}\rightarrow Z\right\rVert^{-\frac{2}{p}}\,\left(\frac{\lambda^{p}-1}{2\lambda^{p}-1}\right)^{\frac{1}{p}}\,&\text{for $p\geq 2$},\\[5.69054pt] \frac{E_{2}(X)}{s\left(tn^{\frac{1}{q}}\right)^{\frac{(2-p)}{p}}}\max\left\{\frac{1}{(\sqrt{n})^{1-\theta}\left\lVert Id:\ell^{n}_{2}\rightarrow Z\right\rVert},\frac{\left\lVert\mathrm{Id}:\ell^{n}_{2}\rightarrow Z\right\rVert^{-\theta}}{e^{(1-\theta)}\left\lVert\mathrm{Id}:Z\rightarrow\ell^{n}_{1}\right\rVert^{1-\theta}}\right\}\left(\frac{\lambda^{p}-1}{2\lambda^{p}-1}\right)^{\frac{1}{p}}\hskip-8.53581pt&\text{for $1<p<2$},\end{cases}$
where $\theta=(2(p-1))/p$, $s:=\left\lVert\mathrm{Id}:Z\rightarrow\ell^{n}_{q}\right\rVert$, and $t:=\left\lVert\mathrm{Id}:\ell^{n}_{q}\rightarrow Z\right\rVert$. Here, $E_{1}(X)$, $E_{2}(X)$, and $E_{3}$ are positive constants: $E_{1}(X)$ depends only on $X$, $E_{2}(X)$ depends on both $X$ and $p$, while $E_{3}$ is independent of $X$ and depends only on $p$. Furthermore,
$L_{\lambda}(B_{Z},p,X)\leq d\,\lambda^{\frac{1}{\log\,n}}\,b_{[\log\,n]}(B_{Z})\,\,\frac{\left\lVert\mathrm{Id}:Z\rightarrow\ell^{n}_{2}\right\rVert}{n^{\frac{1}{p}}}$
@@
Remark 3.1.
The correspondence preceding Theorem 3.1 allows for an extension of Theorem 3.1 to bounded, simply connected, convex complete Reinhardt domains. Furthermore, by combining Theorem 3.1 with Lemma 2.1, we obtain asymptotic estimates for all bounded, simply connected complete Reinhardt domains, not necessarily convex. It is worth noting that Theorem 3.1 may be regarded as a pluriharmonic analogue of [20, Proposition 3.2 and Theorem 3.3], where the authors treated only scalar-valued holomorphic functions with the power $p=1$.
In particular, for the Minkowski spaces $Z=\ell^{n}_{q}$, by proceeding along the same lines as in the proof of Theorem 3.1 and using the estimates obtained in Lemma 3.1, we derive the following bounds. We omit the details.
Corollary 3.1.
Let $X=\mathcal{B}(\mathcal{H})$ be finite dimensional and $\lambda>1$. Then
$L_{\lambda}(B_{\ell^{n}_{q}},p,X)\geq\begin{cases}E^{\prime}_{3}(X)\,\left(\frac{\lambda^{p}-1}{2\lambda^{p}-1}\right)^{\frac{1}{p}}\,\frac{1}{n^{1/q}}\left(\frac{\log n}{n}\right)^{\left(1-\frac{1}{\min\{q,2\}}\right)}&\text{for $p=1$},\\[5.69054pt] E^{\prime}_{2}\left(\frac{\lambda^{p}-1}{2\lambda^{p}-1}\right)^{\frac{1}{p}}\,n^{-\frac{1}{p}}&\text{for $p\geq q$},\\[5.69054pt] E_{4}(X)\,\left(\frac{\lambda^{p}-1}{2\lambda^{p}-1}\right)^{\frac{1}{p}}\,n^{-\left(\frac{p-1}{p(q-1)}+\frac{2-p}{pq}\right)}\,\left(\frac{\log n}{n}\right)^{\left(1-\frac{1}{\min\{q,2\}}\right)\,\frac{q-p}{p(q-1)}}&\text{for $1<p<q$}.\end{cases}$
Here, $E^{\prime}_{3}(X)$, $E^{\prime}_{2}$, and $E_{4}(X)$ are positive constants: $E^{\prime}_{3}(X)$ depends only on $X$, $E_{4}(X)$ depends on both $X$ and $p$, while $E^{\prime}_{2}$ is independent of $X$ and depends only on $p$. Furthermore,
$L_{\lambda}(B_{\ell^{n}_{q}},p,X)\leq d\,\lambda^{\frac{1}{\log\,n}}\,b_{[\log\,n]}(B_{\ell^{n}_{q}})\,\,\frac{\left\lVert\mathrm{Id}:\ell^{n}_{q}\rightarrow\ell^{n}_{2}\right\rVert}{n^{\frac{1}{p}}}$
for some constant $d>0$.
Remark 3.2.
We now highlight two important remarks. First, by arguments analogous to those used in the proof of Theorem 3.1, one can investigate the asymptotic behaviour of the second Bohr radius constant $B_{\lambda}(B_{Z},p,X)$ for holomorphic functions and show that the asymptotic behaviors of $L_{\lambda}(B_{Z},p,X)$ and $B_{\lambda}(B_{Z},p,X)$ coincide up to a multiplicative constant.
Second, we note a difference between the lower bound of $L_{\lambda}(B_{\ell^{n}_{q}},p,X)$ obtained in Corollary 3.1 and the one given in (3.1). In the lower bound appearing in (3.1), no logarithmic term occurs, whereas in our result in Theorem 3.1 the lower bound contains a logarithmic factor. A similar phenomenon was also observed by the author in [18], where only scalar valued holomorphic functions were considered, with $U$ being the identity on $\mathbb{C}$, and with $\lambda=1$ and $p=1$.
@@ 3.1. Basic definitions and notations
$\operatorname{Cot}(W):=\inf\{2\leq t\leq\infty:W\,\mbox{has cotype}\,t\}.$
It is well known that every Banach space has cotype $\infty$. We adopt the convention $1/\operatorname{Cot}(W)=0$ whenever $\operatorname{Cot}(W)=\infty$.
@@ 3.2. Second Bohr radii for homogeneous plynomials
We want to first obtain the desired lower bound of $R^{m}_{\lambda}(B_{Z},p,X)$ when $X$ is finite dimensional. For all $m\geq 1$, let $P(z)=\sum_{|\alpha|=m}a_{\alpha}\,z^{\alpha}+\sum_{|\beta|=m}b^{*}_{\alpha}\,\bar{z}^{\alpha}\in\mathcal{BPH}(^{m}B_{Z},X)$. It is known from [20, Lemma $3.5$] that for $1\leq q\leq\infty$ and $|\alpha|=m$,
$\sup_{z\in B_{\ell_{q}^{n}}}|z^{\alpha}|=\left(\frac{\alpha^{\alpha}}{m^{m}}\right)^{\frac{1}{q}}.$
Now let $B_{Z}$ denote the unit ball of a complex Banach space $Z=(\mathbb{C}^{n},\|\cdot\|)$. We want to find an upper bound for $\sup_{z\in B_{Z}}|z^{\alpha}|$. Since $B_{Z}\subseteq\|\mathrm{Id}:Z\to\ell_{q}^{n}\|\,B_{\ell_{q}^{n}}$, for any $z\in B_{Z}$, we have
$|z^{\alpha}|\leq\sup_{w\in B_{\ell_{q}^{n}}}|w^{\alpha}|\,\|\mathrm{Id}:Z\to\ell_{q}^{n}\|^{\,m},\,\,\,\,\,\,|\alpha|=m.$
Hence,
(3.4) $\sup_{z\in B_{Z}}|z^{\alpha}|\leq\left(\frac{\alpha^{\alpha}}{m^{m}}\right)^{\frac{1}{q}}\,\|\mathrm{Id}:Z\to\ell_{q}^{n}\|^{\,m}$
for $|\alpha|=m$. Note that $\left(\frac{\alpha^{\alpha}}{m^{m}}\right)^{\frac{1}{q}}\leq 1$. Then for $r\in[0,1)$, (3.4) yields
(3.5) $\sum_{|\alpha|=m}\sup_{z\in r\,B_{Z}}(\left\lVert a_{\alpha}\right\rVert^{p}+\left\lVert b_{\alpha}\right\rVert^{p})|z|^{p\alpha}\leq r^{pm}\,\left\lVert\mathrm{Id}:Z\rightarrow\ell^{n}_{q}\right\rVert^{pm}\,\sum_{|\alpha|=m}(\left\lVert a_{\alpha}\right\rVert^{p}+\left\lVert b_{\alpha}\right\rVert^{p}).$
We now need to estimate the term $\sum_{|\alpha|=m}(\left\lVert a_{\alpha}\right\rVert^{p}+\left\lVert b_{\alpha}\right\rVert^{p})$. To do so, we consider several cases depending on the values of $p$. More precisely, the proof is presented in three separate cases, depending on the values of $p$: $p=1$, $p\geq 2$, and $1<p<2$. Let $H$ and $G$ be the $m$-homogeneous polynomials as defined by
$H(z)=\sum_{|\alpha|=m}\frac{a_{\alpha}+b_{\alpha}}{2}\,z^{\alpha},\qquad G(z)=\sum_{|\alpha|=m}\frac{a_{\alpha}-b_{\alpha}}{2}\,z^{\alpha}.$
Set $s:=\left\lVert\mathrm{Id}:Z\rightarrow\ell^{n}_{q}\right\rVert$ and $t:=\left\lVert\mathrm{Id}:\ell^{n}_{q}\rightarrow Z\right\rVert$.
@@
(3.10) $\chi_{M}(\mathcal{P}(^{m}Z))\leq n^{\frac{m}{2}}\,\left\lVert\mathrm{Id}:\ell^{n}_{2}\rightarrow Z\right\rVert^{m}\,\,\mbox{and}\,\,\chi_{M}(\mathcal{P}(^{m}Z))\leq\frac{m^{m}}{m!}\,\left\lVert\mathrm{Id}:Z\rightarrow\ell^{n}_{1}\right\rVert^{m},$
@@
$L^{m}_{\lambda}(B_{Z},p,X)\geq E_{1}(X)\,\,\frac{\lambda^{\frac{1}{m}}}{s\,t\,n^{\frac{1}{q}}}\,\max\left\{\frac{1}{\sqrt{n}\,\left\lVert\mathrm{Id}:\ell^{n}_{2}\rightarrow Z\right\rVert},\,\frac{1}{\left(\frac{m^{m}}{m!}\right)^{1/m}\,\left\lVert\mathrm{Id}:Z\rightarrow\ell^{n}_{1}\right\rVert}\right\},$
@@
$\left(\sum_{|\alpha|=m}\left|\phi\left(\frac{a_{\alpha}+b_{\alpha}}{2}\right)\right|^{2}\right)^{\frac{1}{2}}\leq\left\lVert\mathrm{Id}:\ell^{n}_{2}\rightarrow Z\right\rVert^{m}\left\lVert H\right\rVert_{B_{Z},X}$
@@
(3.11) $\sum_{|\alpha|=m}\left\lVert\left(\frac{a_{\alpha}+b_{\alpha}}{2}\right)\right\rVert^{p}\leq 2^{p-2}\,\left\lVert\mathrm{Id}:\ell^{n}_{2}\rightarrow Z\right\rVert^{2m}\left\lVert H\right\rVert^{2}_{B_{Z},X}.$
@@
(3.12) $\sum_{|\alpha|=m}\left\lVert\left(\frac{a_{\alpha}-b_{\alpha}}{2}\right)\right\rVert^{p}\leq 2^{p-2}\,\left\lVert\mathrm{Id}:\ell^{n}_{2}\rightarrow Z\right\rVert^{2m}\left\lVert G\right\rVert^{2}_{B_{Z},X}.$
@@
$\left(\sum_{|\alpha|=m}\left\lVert a_{\alpha}\right\rVert^{p}\right)^{\frac{1}{p}}\leq 2\left\lVert\mathrm{Id}:\ell^{n}_{2}\rightarrow Z\right\rVert^{\frac{2m}{p}}2^{\frac{2}{p}}.$
@@
$\sum_{|\alpha|=m}\sup_{z\in r\,B_{Z}}(\left\lVert a_{\alpha}\right\rVert^{p}+\left\lVert b_{\alpha}\right\rVert^{p})|z|^{p\alpha}\leq 2^{p+3}\,r^{pm}\,s^{pm}\,\left\lVert\mathrm{Id}:\ell^{n}_{2}\rightarrow Z\right\rVert^{2m}\leq\lambda^{p}$
@@
$r\leq\frac{\lambda^{\frac{1}{m}}}{2^{\frac{p+3}{pm}}\,s\,\left\lVert\mathrm{Id}:\ell^{n}_{2}\rightarrow Z\right\rVert^{\frac{2}{p}}}.$
@@
$L^{m}_{\lambda}(B_{Z},p,X)\geq\frac{\lambda^{\frac{1}{m}}}{2^{\frac{p+3}{pm}}\,s\,\left\lVert\mathrm{Id}:\ell^{n}_{2}\rightarrow Z\right\rVert^{\frac{2}{p}}}.$
@@
$\sum_{|\alpha|=m}\left\lVert\left(\frac{a_{\alpha}+b_{\alpha}}{2}\right)\right\rVert^{p}\leq 2^{p}\,(\pi_{1}(I_{X}))^{2-p}\,\left(t\,n^{\frac{1}{q}}\right)^{(2-p)m}(\chi_{M}(\mathcal{P}(^{m}Z)))^{2-p}\left\lVert\mathrm{Id}:\ell^{n}_{2}\rightarrow Z\right\rVert^{2(p-1)m}.$